# Predicting Forced Responses of Probability Distributions via the Fluctuation–Dissipation Theorem and Generative Modeling

### v1

@@ 3 Results
In what follows, we estimate the mean (first moment) and the second, third, and fourth central moments of the perturbed steady‑state distribution for three representative nonlinear stochastic systems of increasing complexity: a scalar stochastic model for low‑frequency climate variability, a slow‑fast triad model mimicking essential features of the El Niño‑Southern Oscillation (ENSO), and a six‑dimensional stochastic barotropic model capturing atmospheric regime transitions. These systems serve as ideal testbeds for evaluating the accuracy of different approximations to the Generalized Fluctuation-Dissipation Theorem (GFDT) in predicting higher-order responses under small external perturbations.
@@
For each system, we construct the response functions as defined in Eq. (7). Specifically, we focus on impulse response functions for the mean (first moment) and the second, third, and fourth central moments of the invariant distribution by choosing as observables $\bm{x}$, $(\bm{x}-\bm{\mu})^{2}$, $(\bm{x}-\bm{\mu})^{3}$, and $(\bm{x}-\bm{\mu})^{4}$, where $\bm{x}$ denotes the system state and $\bm{\mu}$ its unperturbed mean. Specifically, we assess three distinct approaches: (i) a data-driven method that constructs the response using the score function estimated via the KGMM (K-means Gaussian Mixture Modeling) algorithm; (ii) a Gaussian approximation, which assumes the steady-state distribution to be multivariate Gaussian, fully characterized by its mean and covariance; and (iii) a reference or “ground truth” response. In the case of the scalar triad model, the ground truth has been derived analytically in Majda et al. (2009b) from the known stationary distribution. For the slow-fast triad and stochastic barotropic models, where an analytic form of the score is not available, the ground truth response is computed numerically by performing ensemble integrations of both the unperturbed and perturbed systems and evaluating the change in the moments, see for example the Appendix in Boffetta et al. (2003) or Giorgini et al. (2024b). In all examples, we normalize the data such that the unperturbed time series in each dimension has zero mean and unit variance. This preprocessing step facilitates the comparison across variables and models, and ensures that the Gaussian approximation always corresponds to a standard normal distribution. For systems subject to a state-independent perturbation, as in the first and third examples, we normalize the response functions by the perturbation amplitude, rendering them independent of its magnitude.
@@
In settings where direct computation of response functions is infeasible—either because the underlying equations are unknown or integration is prohibitively expensive—the only reliable way to validate the KGMM‐estimated score (and, by extension, any response predictions) is to compare observed and KGMM‐generated marginal PDFs. Concretely, we simulate the Langevin dynamics
@@
where $\bm{s}(\bm{x})=\nabla\ln\rho_{S}(\bm{x})$ is the score learned via KGMM. The degree to which the simulated marginals reproduce the empirically observed steady‐state PDFs thus provides the sole stringent diagnostic that the score has been learned accurately and that any inferred response functions can be trusted. For this reason, we have also compared the marginal steady‐state PDFs obtained from these Langevin simulations directly against the empirical distributions, offering an independent check on the fidelity of the estimated score function.
@@ 3.1 Scalar Stochastic Model for Low-Frequency Climate Variability
where $\xi(t)$ is standard white noise. This scalar nonlinear model retains the essential deterministic structure of the reduced triad system, while simplifying the stochastic component, thus making it an ideal testbed for analytical and numerical studies.
The coefficients of the model used in this work are listed in Table 1. The input parameters of the KGMM algorithm (see Section 2.2) are set to $\sigma=0.05$ and $N_{C}=348$.
Table 1: Parameters of the stochastic scalar model.
$a$ $-0.0222$
$b$ $-0.2$
$c$ $0.0494$
$F$ $0.6$
$\sigma$ $0.7071$
@@
We now construct impulse response functions through GFDT as $R(t)=\langle A(x(t))\,B(x(0))\rangle_{0}$, and remind the reader that $\langle\cdot\rangle_{0}$ denotes the expectation with respect to the unperturbed steady-state distribution $\rho_{S}(x)$. We consider the simple case where the conjugate observable is defined as $B(x)=-s(x)=-\frac{d}{dx}\log\rho_{S}(x)$. To estimate the score function $s(x)$ we consider three different strategies: (i) analytically from Eq. (24), (ii) using the KGMM method, and (iii) under the Gaussian approximation. In Figure 1, we compare the unperturbed steady-state PDF as well as the first four moment response functions obtained with these three approaches. The results show that the score function estimated via KGMM yields a nearly perfect match with the analytic response, both in terms of the steady‑state PDF and the responses of the mean (first moment) and the second, third, and fourth central moments. This confirms the ability of KGMM to accurately reconstruct the score function. As expected from previous studies (e.g. (Baldovin et al., 2020)), the Gaussian approximation provides a very good estimation for changes in the mean. However, significant biases are present in the second, third and fourth centralized moments, reflecting its inability to capture the non-Gaussian features of the underlying dynamics.
Figure 1: Left: True unperturbed PDF (blue) compared with a normal Gaussian (black) and the PDF obtained integrating Eq. (21) with the KGMM score function (red). Right: Response functions predicted via GFDT. The true response functions (blue) are obtained using the analytic score function and are compared with the ones obtained using the KGMM score function (red) and the linear approximation (black).
@@ 3.2 Slow-Fast Triad Model and Application to ENSO
with parameters listed in Table 2. The input parameters of the KGMM algorithm (see Section 2.2) are set to $\sigma=0.05$ and $N_{C}=7353$.
$d_{u}$ $0.2$
$\omega$ $0.4$
$d_{\tau}$ $2.0$
$\sigma_{u1}=\sigma_{u2}$ $0.3$
$\sigma_{\tau}(u_{1})$ $1.5(\tanh(u_{1})+1)$
Table 2: Parameters of the slow-fast triad model.
@@
In this case, the conjugate observable needs to be considered in its full formulation as $B(\bm{x})=-\nabla\cdot\bm{u}(\bm{x})-\bm{u}(\bm{x})\cdot\nabla\ln\rho_{S}(\bm{x})$. Unlike the triad model, no analytic expression for the steady-state distribution or score function is available for this system. Therefore, the ground truth response was computed by directly simulating an ensemble of both unperturbed and perturbed systems and evaluating the change in moments (see e.g. Giorgini et al. (2024b)). These serve as a benchmark for validating the GFDT-based predictions. As in the previous example, we compared the response of the mean (first moment) and the second, third, and fourth central moments of each variable using three methods: GFDT with the KGMM-estimated score function, GFDT under the Gaussian linear approximation, and the numerical ground truth. We also compared the unperturbed marginal steady-state PDFs predicted by KGMM to the empirically observed distributions by integrating the Langevin equation in Eq. 21.
@@ 3.3 Stochastic Barotropic Model for Atmospheric Regime Transitions
The large-scale circulation of the atmosphere often exhibits persistent patterns, such as zonal and blocked flow regimes, with abrupt transitions between them. These features can be captured through low-order models that isolate the essential mechanisms responsible for regime persistence and transitions (Charney and DeVore, 1979). The six-dimensional model considered here is based on a Galerkin truncation of the barotropic vorticity equation on a $\beta$-plane channel with topography. The original formulation was introduced by Charney and DeVore (Charney and DeVore, 1979), and the specific version used here follows a slightly modified form proposed by De Swart (De Swart, 1988). This model has been widely studied, particularly in the context of stochastic regime dynamics (e.g., Crommelin et al., 2004; Dorrington and Palmer, 2023). It retains two zonal modes and two pairs of wave modes, enabling nonlinear interactions between the waves and the mean flow, as well as capturing the influence of topographic forcing. Physically, the model describes the evolution of the barotropic streamfunction field, subject to planetary rotation, Newtonian damping toward a prescribed zonal background state, and orographic forcing. The six prognostic variables represent amplitudes of selected Fourier modes, and the equations contain linear terms corresponding to damping and rotation, as well as quadratic nonlinearities encoding advective interactions. The system supports multiple equilibria and intermittent transitions between weakly chaotic and more persistent quasi-stationary states. By adding stochastic forcing, one can probe how unresolved processes affect the stability and persistence of regimes (Dorrington and Palmer, 2023). The stochastic version of the model takes the form:
@@
The values of the coefficients used in this study are summarized in Table 3, corresponding to a system with channel aspect ratio $b=1.6$, $\beta=1.25$, and topographic forcing strength $\gamma=0.2$. The forcing terms $x_{1}^{*}=0.95$ and $x_{4}^{*}=-0.76095$ define the equilibrium zonal profile. The input parameters of the KGMM algorithm (see Section 2.2) are set to $\sigma=0.05$ and $N_{C}=30309$.
Parameter Value Description
$\alpha_{1}$ 0.86322463 Nonlinear advection (mode 1)
$\alpha_{2}$ 0.81394451 Nonlinear advection (mode 2)
$\beta_{1}$ 0.89887640 Coriolis effect (mode 1)
$\beta_{2}$ 0.48780488 Coriolis effect (mode 2)
$\delta_{1}$ 1.38115941 Triad interaction (mode 1)
$\delta_{2}$ $-0.12882575$ Triad interaction (mode 2)
$\tilde{\gamma}_{1}$ 0.19206748 Orographic forcing (mode 1)
$\tilde{\gamma}_{2}$ 0.07682699 Orographic forcing (mode 2)
$\gamma_{1}$ 0.05395154 Orographic damping (mode 1)
$\gamma_{2}$ 0.04684573 Orographic damping (mode 2)
$\varepsilon$ 1.44050611 Wave-wave interaction
$C$ 0.1 Newtonian relaxation rate
$x_{1}^{*}$ 0.95 Zonal background forcing (mode 1)
$x_{4}^{*}$ $-0.76095$ Zonal background forcing (mode 2)
$\sigma$ 0.01 Noise amplitude
@@ 4 Discussion: Limitations and practical considerations
Our results show that when (i) a large amount of data is available and (ii) the full state vector $\bm{x}(t)$ is known, the system’s probability density response to small external perturbations can indeed be recovered from the Generalized Fluctuation-Dissipation Theorem (GFDT), as expected from theoretical arguments (Castiglione et al., 2008). Specifically, response functions can be estimated with high precision by leveraging recent advances in score-based generative modeling, such as the KGMM algorithm. The experiments presented here represent a valuable first step toward an equation-free framework for studying how the probability distribution of real-world dynamical systems responds to small external perturbations.
@@
However, several limitations and caveats must be considered before applying these tools to high-dimensional complex systems. A straightforward practical limitation arises when only relatively short time series are available. As with many machine learning-based approaches, the performance of our method depends heavily on the sample size, which in turn is determined by the complexity and temporal scales of the phenomena under study. Even with access to infinitely long time series, a more fundamental issue remains in data-driven approaches: in most real-world scenarios, we rarely observe (or even know) the full state vector of the system. Instead, we typically have access to only a small subset of variables—a projection of the full dynamics. This longstanding problem, emphasized at least since Onsager and Machlup (1953) (Onsager and Machlup, 1953), remains a central concern in modern studies (Cecconi et al., 2012; Hosni and Vulpiani, 2018; Baldovin et al., 2018, 2020). One might be tempted to address this challenge using the Takens embedding theorem (Takens, 1981), reconstructing the underlying dynamics from partial observations. However, this strategy is severely limited in practice: it is not applicable to stochastic systems and becomes rapidly infeasible as the system dimensionality increases; see Cecconi et al. (2012); Baldovin et al. (2018); Lucente et al. (2022) for an in-depth discussion. For these reasons, Takens-based approaches are not a reliable option for studying many real-world systems.
A more promising approach—aligned with the perspective taken in this work and with a long history in climate science—is to exploit the multiscale structure of real-world dynamical systems and focus on coarse-grained effective dynamics. In practice, in certain cases, it is possible to identify slow variables that exhibit low-dimensional effective dynamics. The cumulative effect of unresolved fast variables on the slow ones can then be modeled as stochastic forcing terms (Hasselmann, 1976; Penland, 1989; Majda et al., 2008c; Lucarini and Chekroun, 2023). Response theory can then be applied by focusing on the coarse-grained dynamics (Lacorata and Vulpiani, 2007). Accordingly, the success of the proposed method in real-world systems depends critically on the choice of the proper variables used to study the phenomena of interest; see also Appendix A in Falasca et al. (2025). In practice, these choices are informed by physical intuition and prior knowledge, and they are shaped by the specific coarse-graining and data preprocessing procedures applied. Consequently, a purely “black box” approach is inadequate: the inference of responses functions in complex systems through data-driven methods requires deep physical insight and strong domain expertise.

### v2

@@ 3 Results
In what follows, we estimate the mean (first moment) and the second, third, and fourth central moments of the perturbed steady‑state distribution for four representative nonlinear stochastic systems of increasing complexity: three reduced-order models (a scalar stochastic model for low‑frequency climate variability, a slow‑fast triad model for ENSO, and a six‑dimensional barotropic model for atmospheric regimes), and one high-dimensional partial differential equation (a 2D Navier-Stokes model for turbulent flow). These systems serve as ideal testbeds for evaluating the accuracy of different approximations to the Generalized Fluctuation-Dissipation Theorem (GFDT) in predicting higher-order responses under small external perturbations.
@@
For each system, we construct the response functions as defined in  (7). Specifically, we focus on response functions for the mean (first moment) and the second, third, and fourth central moments of the invariant distribution by choosing as observables $\bm{x}$, $(\bm{x}-\bm{\mu})^{2}$, $(\bm{x}-\bm{\mu})^{3}$, and $(\bm{x}-\bm{\mu})^{4}$, where $\bm{x}$ denotes the system state and $\bm{\mu}$ its unperturbed mean. Additionally, for the 2D Navier–Stokes example, we also analyze the fifth central moment $(\bm{x}-\bm{\mu})^{5}$. We assess three distinct approaches: (i) a data-driven method that constructs the response using the score function estimated via either the KGMM algorithm for the low-dimensional models or a U-Net for the high-dimensional system; (ii) a Gaussian approximation, which assumes the steady-state distribution to be multivariate Gaussian, fully characterized by its mean and covariance; and (iii) a reference or “ground truth” response. In the case of the scalar model, the ground truth has been derived analytically in Majda et al. (2009b) from the known stationary distribution. For the other models, where an analytic form of the score is not available, the ground truth response is computed numerically by performing ensemble integrations of both the unperturbed and perturbed systems and evaluating the change in the moments, see for example the Giorgini et al. (2024b); Falasca (2025) as well as the Appendix of Boffetta et al. (2003). In all examples, we normalize the data such that the unperturbed time series in each dimension has zero mean and unit variance. This preprocessing step facilitates the comparison across variables and models, and ensures that the Gaussian approximation always corresponds to a standard normal distribution. For systems subject to a state-independent perturbation, as in the first and third examples, we normalize the response functions by the perturbation amplitude, rendering them independent of its magnitude.
@@
In settings where direct computation of response functions is infeasible—either because the underlying equations are unknown or integration is prohibitively expensive—the only reliable way to validate the estimated score (and, by extension, any response predictions) is to compare observed and generative model-based marginal PDFs. Concretely, we simulate the Langevin dynamics
@@
where $\bm{s}(\bm{x})=\nabla\ln\rho_{S}(\bm{x})$ is the score learned via our data-driven methods. The degree to which the simulated marginals reproduce the empirically observed steady‐state PDFs thus provides the sole stringent diagnostic that the score has been learned accurately and that any inferred response functions can be trusted. For this reason, we have also compared the marginal steady‐state PDFs obtained from these Langevin simulations directly against the empirical distributions, offering an independent check on the fidelity of the estimated score function.
@@ 3.1 Scalar Stochastic Model for Low-Frequency Climate Variability
where $\xi(t)$ is standard white noise. These dynamics can be interpreted as the motion of an overdamped particle in a one-dimensional potential energy landscape $U(x)$, where the deterministic force is given by F(x)=dU/dxF(x)=-dU/dx. The potential is a quartic polynomial of the form $U(x)=-(Fx+\frac{a}{2}x^{2}+\frac{b}{3}x^{3}-\frac{c}{4}x^{4})$.
The coefficients of the model used in this work are listed in Table 1 of Appendix § 5. The input parameters of the KGMM algorithm (see Section 2.2) are set to $\sigma_{G}=0.05$ and $N_{C}=348$.
@@
We now construct impulse response functions through GFDT as $R(t)=\langle A(x(t))\,B(x(0))\rangle_{0}$, and remind the reader that $\langle\cdot\rangle_{0}$ denotes the expectation with respect to the unperturbed steady-state distribution $\rho_{S}(x)$. We consider the simple case where the conjugate observable is defined as $B(x)=-s(x)=-\frac{d}{dx}\log\rho_{S}(x)$. To estimate the score function $s(x)$ we consider three different strategies: (i) analytically from  (18), (ii) using the KGMM method, and (iii) under the Gaussian approximation. In Figure 1, we compare the unperturbed steady-state PDF as well as the first four moment response functions obtained with these three approaches. The results show that the score function estimated via KGMM yields a nearly perfect match with the analytic response, both in terms of the steady‑state PDF and the responses of the mean (first moment) and the second, third, and fourth central moments. This confirms the ability of KGMM to accurately reconstruct the score function. As expected from previous studies (e.g. (Baldovin et al., 2020)), the Gaussian approximation provides a very good estimation for changes in the mean. For the higher-order moments, it can sometimes capture the qualitative nature of the response, such as the initial sign and relaxation timescale for the variance. However, it introduces significant quantitative biases and becomes increasingly unreliable for higher orders, reflecting its inability to capture the non-Gaussian features of the underlying dynamics.
Figure 1: Left: True unperturbed PDF (blue) compared with a normal Gaussian (black) and the PDF obtained integrating  (15) with the KGMM score function (red). Right: Response functions predicted via GFDT. The true response functions (blue) are obtained using the analytic score function and are compared with the ones obtained using the KGMM score function (red) and the linear approximation (black). Note that the non-zero response of the second moment under the Gaussian approximation is due to the quasi-Gaussian framework used (see Section 2.1), where the expectation is taken over the true non-Gaussian distribution.
@@ 3.2 Slow-Fast Triad Model and Application to ENSO
with parameters listed in Table 2 of Appendix § 5. The input parameters of the KGMM algorithm (see Section 2.2) are set to $\sigma_{G}=0.05$ and $N_{C}=7353$.
@@
In this case, the conjugate observable needs to be considered in its full formulation as $B(\bm{x})=-\nabla\cdot\bm{u}(\bm{x})-\bm{u}(\bm{x})\cdot\nabla\ln\rho_{S}(\bm{x})$. Unlike the triad model, no analytic expression for the steady-state distribution or score function is available for this system. Therefore, the ground truth response was computed by directly simulating an ensemble of both unperturbed and perturbed systems and evaluating the change in moments (see e.g. Giorgini et al. (2024b)). These serve as a benchmark for validating the GFDT-based predictions. As in the previous example, we compared the response of the mean (first moment) and the second, third, and fourth central moments of each variable using three methods: GFDT with the KGMM-estimated score function, GFDT under the Gaussian linear approximation, and the numerical ground truth. We also compared the unperturbed marginal steady-state PDFs predicted by KGMM to the empirically observed distributions by integrating the Langevin equation in (15).
@@ 3.3 Stochastic Barotropic Model for Atmospheric Regime Transitions
The large-scale circulation of the atmosphere often exhibits persistent patterns, such as zonal and blocked flow regimes, with abrupt transitions between them. These features can be captured through low-order models that isolate the essential mechanisms responsible for regime persistence and transitions (Charney and DeVore, 1979). The six-dimensional model considered here is based on a Galerkin truncation of the barotropic vorticity equation on a $\beta$-plane channel with topography. The original formulation was introduced by Charney and DeVore (Charney and DeVore, 1979), and the specific version used here follows a slightly modified form proposed by De Swart (De Swart, 1988). This model has been widely studied, particularly in the context of stochastic regime dynamics (e.g., Crommelin et al., 2004; Dorrington and Palmer, 2023; Grafke and Vanden-Eijnden, 2019). It retains two zonal modes and two pairs of wave modes, enabling nonlinear interactions between the waves and the mean flow, as well as capturing the influence of topographic forcing. Physically, the model describes the evolution of the barotropic streamfunction field, subject to planetary rotation, Newtonian damping toward a prescribed zonal background state, and orographic forcing. The six prognostic variables represent amplitudes of selected Fourier modes, and the equations contain linear terms corresponding to damping and rotation, as well as quadratic nonlinearities encoding advective interactions. The system supports multiple equilibria and intermittent transitions between weakly chaotic and more persistent quasi-stationary states. By adding stochastic forcing, one can probe how unresolved processes affect the stability and persistence of regimes (Dorrington and Palmer, 2023). The stochastic version of the model takes the form:
@@
The values of the coefficients used in this study are summarized in Table 3 of Appendix § 5, corresponding to a system with channel aspect ratio $b=1.6$, $\beta=1.25$, and topographic forcing strength $\gamma=0.2$. The forcing terms $x_{1}^{*}=0.95$ and $x_{4}^{*}=-0.76095$ define the equilibrium zonal profile. The input parameters of the KGMM algorithm (see Section 2.2) are set to $\sigma_{G}=0.05$ and $N_{C}=30309$.
@@
3.4 Two-dimensional Navier-Stokes equation
As a representative example of a high-dimensional system where reduced-order modeling is not applicable, we consider the two-dimensional Navier-Stokes model as described in (Flierl and Souza, 2024; Giorgini et al., 2024b), where we take the state variable to be the vorticity $\zeta$. This system was chosen as a stepping-stone towards the more complex fluids observed in climate systems. The data-generating equation is:
$\partial_{t}\zeta$ $=\partial_{x}\psi\partial_{y}\zeta-\partial_{x}\zeta\partial_{y}\psi-A\partial_{x}\zeta+\mathcal{D}\zeta+\varsigma$ (23)
$\psi$ $=\Delta^{-1}\zeta$ (24)
$\mathcal{D}\zeta$ $=-\nu_{h}\Delta^{-2}\zeta-\nu\Delta^{2}\zeta-\nu_{0}\int\zeta\,\mathrm{d}x\,\mathrm{d}y$ (25)
where $\mathcal{D}$ is a dissipation operator with constants $\nu_{h}=10^{-2}$, $\nu=10^{-5}$, $\nu_{0}=(2\pi)^{-2}$, $A=\pi$ is a mean advection term, and $\varsigma$ is a random wave forcing. The random wave forcing introduces hidden dynamics into the system, thus the score function is estimated using only partial state information. The inverse Laplacian is defined to preserve the mean of the original variable.
The system is discretized on a $32\times 32$ periodic grid using a pseudospectral method. Thus the system consists of $1024$ degrees of freedom in the Navier-Stokes equations. The forcing term maintains energy injection at large scales while viscous dissipation removes energy at small scales, establishing a turbulent steady state. The stochastic forcing $\varsigma$ represents small-scale fluctuations and ensures ergodicity of the system.
We consider a localized perturbation applied to the vorticity field at a single pixel in the computational domain, representing the most localized forcing possible in the discretized system. This single-pixel perturbation couples with all spatial modes of the system, preventing the use of reduced-order models for response prediction and representing an extreme case where external forcing is applied at a single spatial location.
For this high-dimensional system, we collect $N=10^{5}$ samples from the unperturbed steady state and train the U-Net score estimator described above. The network is trained for 500 epochs with a batch size of 64 using the Adam optimizer. The observables we considered were the moments of the probability density function. Due to the reflection symmetry of the vorticity field in this system (where positive and negative vorticity are equally probable), the even moments exhibit zero response to the localized perturbation, as the perturbation preserves this antisymmetric property of the underlying dynamics.
First, we note that the Gaussian approximation provides a very good, first-order estimation of the linear response in the 1st and 3rd moment. This further underscores the utility of the Gaussian approximation as a baseline for more complex methodologies. Second, our results clearly demonstrate the benefit of generative modeling to estimate linear responses in higher-order moments through the GFDT and the methodology in Giorgini et al. (2024b). Figure 4 demonstrates that the U-Net approach provides significantly better agreement with the numerical ground truth compared to the linear approximation, particularly for the higher-order moments where non-Gaussian effects become more pronounced. This example demonstrates the capability of the U-Net-based approach to handle complex, high-dimensional systems where traditional reduced-order modeling fails due to the spatially localized nature of the perturbation and the absence of clear scale separation in the turbulent flow.
Figure 4: Comparison of response functions for the 2D Navier-Stokes system with single-pixel perturbation. The plots show the response of the 1st, 3rd, and 5th moments across different time lags for five different observables, calculated using numerical integration (blue), U-Net score estimation (red), and linear Gaussian approximation (black). The U-Net method demonstrates superior agreement with the numerical ground truth compared to the linear approximation, particularly for higher-order moments where non-Gaussian effects are more significant.
@@ 4 Discussion: Limitations and practical considerations
Our results show that when (i) a large amount of data is available and (ii) the full state vector $\bm{x}(t)$ is known, the system’s probability density response to small external perturbations can indeed be recovered from the Generalized Fluctuation-Dissipation Theorem (GFDT), as expected from theoretical arguments (Castiglione et al., 2008). Specifically, we have demonstrated two complementary data-driven strategies whose selection hinges on the intrinsic spatial organization of the data: for systems that are effectively lower-dimensional in state space (our scalar, triad, and barotropic ROMs), we estimate the score with KGMM and a multilayer perceptron; when the spatial organization is higher-dimensional (e.g., the 2D Navier–Stokes vorticity field on a $32\times 32$ grid), we estimate the score with denosing score-matching and a U-Net. See Appendix § 6.3 for guidance and a side-by-side comparison, and Appendix § 6.1 and Appendix § 6.2 for method details. The experiments presented here represent a valuable step toward an equation-free framework for studying how the probability distribution of real-world dynamical systems responds to small external perturbations.
However, several limitations and caveats must be considered when applying these tools to high-dimensional, “real-world” complex systems. We discuss such caveats below:
[…]
@@
  • Partially observed state vector. Even with access to infinitely long time series, a more fundamental issue remains: in most real-world scenarios, we rarely observe (or even know) the full state vector of the system. Instead, we typically have access to only a small subset of variables—a projection of the full dynamics. This longstanding problem, emphasized at least since Onsager and Machlup (1953) (Onsager and Machlup, 1953), remains a central concern in modern studies (Cecconi et al., 2012; Hosni and Vulpiani, 2018; Baldovin et al., 2018, 2020). One might be tempted to address this challenge using the Takens embedding theorem (Takens, 1981), reconstructing the underlying dynamics from partial observations. However, this strategy is severely limited in practice: it is not applicable to stochastic systems and becomes rapidly infeasible as the system dimensionality increases; see Cecconi et al. (2012); Baldovin et al. (2018); Lucente et al. (2022) for an in-depth discussion. For these reasons, Takens-based approaches are not a reliable option for studying many real-world systems. The reduced order modeling approach, as in the first part of this paper, attempts to tackle this fundamental limitation, by exploiting the multiscale structure of real-world dynamical systems and focusing on coarse-grained effective dynamics. The success of GFDT in this context hinges on two key conditions: a clear timescale separation between the resolved slow modes and unresolved fast modes, and a perturbation that projects primarily onto these slow modes. When these conditions hold, as is common for large-scale climate forcings, the cumulative effect of unresolved variables can be modeled as stochastic noise, and response theory can be effectively applied to the coarse-grained dynamics (Hasselmann, 1976; Penland, 1989; Majda et al., 2008c; Lucarini and Chekroun, 2023; Lacorata and Vulpiani, 2007; Lucarini, 2025; Falasca, 2025). The success of this proposed method in real-world systems depends critically on the choice of the proper variables used to study the phenomena of interest; we refer the reader to the contribution in Falasca (2025) for more details. While our results focused on data-driven strategies applied to pre-defined ROMs, future work will extend these ideas to the more challenging scenario of inferring responses directly from partially observed data of high-dimensional systems.
  • Absence of scale separation. Finally, it is possible to envision conditions where we want to study the response to perturbations coupling with all modes. The feasibility of such approach will depend on observing of the full state vector, or at least all the relevant proper variables. Our analysis of the 2D Navier-Stokes system, with its localized perturbation coupling to all spatial scales, exemplifies such a case. Here, a ROM is inadequate, and a direct estimation of the score function on the high-dimensional state space is required. Our use of a U-Net architecture demonstrates that this is a viable, albeit more computationally demanding, alternative.